OpenStudy (anonymous):

Solve by the elimination method 3r-2s=19 2r+3s=30

OpenStudy (anonymous):

29

OpenStudy (anonymous):

hello diana

OpenStudy (diana):

-2(3r-2s=19) --> -6r+4s=-38 3(2r+3s=30) --> 6r+9s=90 Can you solve the rest?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i always have trouble with these

OpenStudy (anonymous):

5s=128

OpenStudy (anonymous):

divide 128divided by 5s

OpenStudy (anonymous):

25.6

OpenStudy (diana):

wait where did you get 5s? its 13s=52 then you divide by 13 and get s=4

OpenStudy (anonymous):

2s+3s

OpenStudy (diana):

oh. okay. remember how i multiplied one equation by -2 and the other by 3, well that gives you -6r+4s=-38 and 6r+9s=90. these are your new equations and it allows you to cancel out r and solve for s.

OpenStudy (anonymous):

So the solution is what

OpenStudy (diana):

s=4 and r=9

OpenStudy (diana):

do you want me to show you the steps?

OpenStudy (anonymous):

yes

OpenStudy (diana):

s=4 and r=9

OpenStudy (diana):

sorry hold on

OpenStudy (diana):

step one: look for a number that can you can multiply each equation to make one of the coefficients in front of the variable cancel out... -2(3r-2s=19) --> -6r+4s=-38 3(2r+3s=30) --> 6r+9s=90 step 2: when adding the two new equations -6r and 6r cancels out while 4s+9s=13s and -38+90=52. this gives you 13s=52 step 3: divide by 13 to both sides and s=4 step 4: plug in s into one of the original equations and you solve for r soo.. 3r-2(4)=19 3r-8=19 step 5: add 8 to both sides 3r=27 step 6: divide by 3 on both sides r=9 there you go. hope that helps.

OpenStudy (anonymous):

ty